Find the perimeter of a rectangle if three of its corners are (2.5,1.7), (2.5,-3.2), and (-0.7,-3.2).
Answer by venugopalramana(3286)

Find the perimeter of a rectangle if three of its corners are A= (2.5,1.7), B=(2.5,-3.2), and C= (-0.7,-3.2).
AB=SQRT{(2.5-2.5)^2+(1.7+3.2)^2}=4.9
BC=SQRT{(2.5+0.7)^2+(-3.2+3.2)^2}=3.2
HENCE PERIMETER=2(4.9+3.2)=16.2